OpenStudy (lukecrayonz):

The decay rate, k, for a particular radioactive element is 2.8%, where time is measured in years. Find the half-life of the element

OpenStudy (lukecrayonz):

@timo86m :D

OpenStudy (lukecrayonz):

@ranga

OpenStudy (ranga):

Half-life = ln(2) / k

OpenStudy (lukecrayonz):

ln(2)/0.028?

OpenStudy (ranga):

Half life = ln(2) / .028 = 24.8 years

OpenStudy (lukecrayonz):

almost used log in my answer, got 34 :P thank you.

OpenStudy (ranga):

yw.

OpenStudy (wolf1728):

half life = ln(2) / decay constant half life = .693 / .028 half life = 24.75 years